### v2

# theoTheorem

Abstract
Laser phase noise is one of the main sources of decoherence in driven Rydberg systems with neutral atoms in tweezer arrays. While the effect of phase noise in the regimes of isolated qubits and few-qubit gate protocols has been studied extensively, there are open questions about the effects of this noise on many-body systems. In many scenarios, the effects of noise cannot simply be described by an increase in the energy or temperature of the system, leading to non-trivial changes in the state and relevant correlations. In this work, we use stochastic sampling to simulate laser phase noise based on experimentally relevant spectral densities. We explore the impact of this noise on adiabatic state preparation in a one-dimensional system, discussing the interplay between heating and interactions during dynamics. We find that for certain adiabatic processes, the noise can be seen to approximately thermalize in terms of its effects on relevant correlation functions.
preprint: APS/123-QED
I Introduction
Progress on the scalability of optical trapping and coherent control have made neutral atom arrays an exciting platform for quantum computation and simulation [1, 2, 3, 4]. In these systems, strong inter-particle interactions between individually trapped atoms are realized by exciting atoms to high-energy Rydberg levels, giving rise to a blockade effect [5] that can be leveraged for the creation of quantum gates [6, 7, 8], as well as to a range of spin Hamiltonians [9, 34]. Applications of these systems include the exploration of annealing processes relevant for optimization [10, 11], realizing topological models [12, 13], spin liquid phases [14, 15, 16], quantum many-body scars [17] and information scrambling [18, 19, 20]. Much of the progress in this area in the past years came from reductions in laser phase noise, which crucially affects excitations to Rydberg levels and directly generates dephasing [21, 22, 23, 24]. The effects of this laser phase noise have mainly been quantified in the context of single qubit operations. In simulation of quantum dynamics, an important question is how this single-particle heating affects the overall behavior of a many-body system. This connects to fundamental questions on thermalization in closed quantum systems [25, 26, 27, 28, 29], specifically whether the heating due to laser phase noise results in correlation functions that would be predicted by the increase in energy and temperature.
In this work, we quantify the effects of laser phase noise on a many-body spin systems as realized in tweezer arrays. The adiabatic preparation of many-body ground states provides an important example in this context, with the procedure used to prepare interesting many-body states [30, 31, 32, 9], as well as in quantum annealing [33, 34, 35, 36]. We simulate time evolution of an adiabatic state preparation of our antiferromagnetic spin chain, applying laser phase noise that has been sampled from experimentally realistic diode laser data, similar to phase noise measured, e.g., in Ref. [21]. We extract features that are relevant to general noise spectra, quantifying the reduction of fidelity and changes in correlation functions that result from adding phase noise with varying strengths and frequency profiles relative to central frequencies associated with the system dynamics. We analyze in detail the couplings to excited states generated by noise, and also show to what extent (despite the complicated many-body spectra) sensitivity to noise can be reduced through the choice of the relative frequencies. We then also ask to what extent we might have expected the results we obtain for correlations based on thermalization of the states. We evaluate the long-time average expectation values of relevant observables, comparing them to predictions by a thermal equilibrium state of the same energy to see how closely the relaxation of energy in the system approaches thermalization.
The manuscript is organized as follows: after an introduction to the physical system and model in Section II, Section III considers the effects of phase noise on the simulations of an adiabatic state preparation of a one-dimensional Rydberg spin chain with antiferromagnetic ($Z_{2}$) ordering. Section IV introduces the concept of thermalization, followed by a demonstration of thermalization in two separate many body observables. Finally, Section V provides a concluding discussion on the analysis of laser phase noise in Rydberg systems conducted in this paper. We contextualize the findings within the broader advances in state-of-the-art neutral atom systems, highlighting how insights gained from this work underscore the importance of understanding and mitigating phase noise effects in many-body quantum systems.
Figure 1: Parameter space plot for the Rydberg Hamiltonian seen in Eq. (II) for a chain of N=11 atoms. (a) A schematic diagram showing crystalline ordering of ground (blue) and Rydberg excited (red) atoms for initial and target many-body state of adiabatic state preparation. (b) Total fraction of excited Rydberg atoms $N_{R}/N$ in the ground state plotted against laser parameters $\Omega/\delta$, from no excitations (blue), to half filled (white), to fully excited (red).The path for the diabatic state preparation studied is provided with color gradient from white to green tracking ground state energy gap. All parameters are provided in units interaction strength $V_{dd}$.
II The physical system
The simulations discussed in this paper are carried out in a one-dimensional Rydberg spin Hamiltonian that takes the form
$\hat{H}_{R}=\quad$ $\frac{\Omega(t)}{2}\sum^{N}_{k}\big{[}e^{-i\phi(t)}|0\rangle_{k}\langle 1|_{k}+e^{i\phi(t)}|1\rangle_{k}\langle 0|_{k}]$
$-\delta(t)\sum^{N}_{k}\hat{n}_{k}+V_{dd}\sum^{N}_{k,l>k}\frac{\hat{n}_{k}\hat{n}_{l}}{|k-l|^{3}},$ (1)
where the first term drives excitations in the system at a Rabi frequency $\Omega(t)$ modulated by a phase noise signal $\phi(t)$ that introduces a small error at every time step. The parameter $\delta(t)$ models the global laser detuning, and the operator $n_{k}\equiv\sigma_{z}^{k}-\hat{I}$ counts the number of Rydberg excitations in site $k$. The terms $\delta(t)$ and $\Omega(t)$ are tuneable in an experimental setting and are governed by the laser frequency, amplitude, and phase stability. Finally, we define the interaction Hamiltonian $\hat{H}_{\text{int}}$ explicitly to use it as an observable
$\hat{H}_{\text{int}}=V_{dd}\sum^{N}_{k,l>k}\frac{\hat{n}_{k}\hat{n}_{l}}{|k-l|^{3}},$ (2)
introducing a dipole-dipole potential $V_{dd}$ mediated by $\hat{n}_{k}\hat{n}_{l}$ interactions that enforce a Rydberg blockade penalty on neighboring excitations. We work in the dipole dipole regime in which long range interactions decay as a function of distance between sites $d^{3}=|k-l|^{3}$, where $k$ and $l$ are site indices and the lattice constant is absorbed into $V_{dd}$. Open boundary conditions are considered for the one-dimensional chain. For computational simplicity, we define a relative number to which we set the interaction strength $V_{dd}/2\pi=1/2\pi$  MHz, and use this to scale all units in the system with $\hbar=1$ also assumed. Hence, both $\Omega$ and $\delta$ are expressed in units of this relative $V_{dd}$, while time is given in units of $V_{dd}^{-1}$. The choice allows us to study many body dynamics in a similar frequency region as the sampled noise, and the results in this paper can be generalized to different frequency regimes. Note that for $\Omega=0$ the Hamiltonian in Eq. (II) is reminiscent of the Ising model for magnetic dipoles. In such a regime, the competition between $\delta$ and $V_{dd}$ leads to an excitation ladder in which ground states with increasing number of evenly spaced excitations are favored [30, 32, 37]. These crystalline ground states exhibit large energy gaps that make them robust to small increases in $\Omega$, leading to regions in the $\Omega/\delta$ parameter space where only one crystalline state is represented in the ground state. In Fig. 1 we show Rydberg excitations along the spin chain as a function of $\Omega/\delta$ parameters, clearly displaying the formation of a lobe with a stable ground state with crystalline $Z_{2}$ order of excitations. How close the final prepared many-body ground state $|\psi_{gr}\rangle$ is to a perfect $Z_{2}$ ordering can always be tested by computing the expectation value of the order parameter observable
$\hat{O}_{Z_{2}}=\sum^{N}_{k,l\neq k}\sigma_{z}^{k}(-1)^{k+l}\sigma_{z}^{l},$ (3)
where we check anti-ferromagnetism over all different lattice sites $k$ and $l$. The goal of the procedure is to take a ground state with no Rydberg excitations to a ground state with $Z_{2}$ ordered excitations and expectation value $\langle\hat{O}_{Z_{2}}\rangle=1$, while maximizing the ground state energy gap to mitigate diabatic excitation. Due to the open boundary conditions, atoms found on the edges of the chain experience less of the effect of Rydberg blockade and are therefore more likely to be Rydberg excited than atoms within the chain. This means that energetically favorable $Z_{2}$ ordering of excitations can be identified uniquely in chains with an odd number of lattice sites, but an even number of lattice sites results in a central domain wall after which ordering flips [9]. To simplify notation and discussions, here we will focus on odd numbers of spins. Also shown in Fig. 1 is the path of the adiabatic state preparation simulated in this paper, broken up into a series of three linear ramps that tune $\Omega$ and $\delta$ independently. The next section will introduce this state preparation in more detail as well as providing results after introducing phase noise to the system.
III Phase noise in adiabatic state preparation
Figure 2: (a) Laser parameters for three step adiabatic state preparation of a $Z_{2}$ ordered crystalline state for a 11 atom Rydberg chain. Rabi frequency $\Omega$ (red) is tuned linearly in step 1,3 for periods $T_{1}$, $T_{3}$, while detuning $\delta$ (green) is tuned linearly only in step 2 for a period of $T_{2}$. (b) Concurrent fidelity measured as the wavefunction overlap between the ground state $|\psi_{gr}\rangle$ of the instantaneous Hamiltonian $\hat{H}_{R}(t)$ and the evolved state $|\psi\rangle$. (c) Energy gap $\Delta E_{gr}$ between the instantaneous ground state and the first excited state.
Adiabatic state preparation is a technique to initialize a quantum system in a desired state [37, 31]. The method is useful for preparing complex quantum states that are otherwise difficult to generate directly. It typically involves slowly varying the controllable parameters of a system to transform a given energy eigenstate state in an initial Hamiltonian to a target state that is the ground state of the modified Hamiltonian. This variation should be slow enough to satisfy the adiabatic constraint
$|\langle E_{m}(t)|\dot{H}(t)|E_{n}(t)\rangle|\ll|E_{m}(t)-E_{n}(t)|,$ (4)
where $|E_{m}(t)\rangle$ and $|E_{n}(t)\rangle$ are eigenstates with energies $E_{m}(t)$ and $E_{n}(t)$ in the instantaneous Hamiltonian with an energy gap that we do not wish to cross. Typically, the procedure is done starting from a state which is easy to prepare experimentally to a target state which is harder to access. The challenge in adiabatic state preparation is thus to maximize the energy gap to the first excited state $\Delta E_{gr}$ and to decrease the variation rate if $\Delta E_{gr}$ becomes small enough to result in unwanted excitation of higher energy eigenstates.
III.1 The adiabatic ramp
Figure 3: Realistic noise power spectrum used to generate independent phase noise time realizations implemented in adiabatic time evolution. (a) shows the phase noise spectrum $S_{\phi}$ (green) which has been converted from a frequency spectrum $S_{\nu}$ (red) used in noise generation, with dashed line marking system interaction strength $V_{dd}$. The relevant conversion equation is $S_{\phi}=\nu^{2}S_{\nu}$. Also provided in (b) are three independent noise realizations generated using $S_{\phi}$ (see Appendix A). The averaged noise behavior after 100 independently generated noise realizations is shown in blue.
Both laser parameters $\delta$ and $\Omega$ in the Hamiltonian in Eq. (II) are time dependent and can be tuned independently. Starting from a Hamiltonian with large negative detuning in a ground state with no Rydberg excitations, Fig. 2(a) shows the preparation of the $Z_{2}$ target state using a three-step adiabatic ramp originally proposed in [31]. The initial Hamiltonian is far detuned at laser parameters $\Omega/V_{dd}=0.1$ and $\delta/V_{dd}=-3$ where a large ground state gap allows for easy experimental preparation. The target Hamiltonian occurs at $\Omega/V_{dd}=0.1$ and $\delta/V_{dd}=1.1$ with a much smaller gap and a ground state with $Z_{2}$ excitation ordering. Figures 2(b,c) show the change in the ground state fidelity $|\langle\psi|\psi_{gr}\rangle|^{2}$ and ground state energy gap $\Delta E_{gr}$ of such an adiabatic state preparation for a chain of 11 atoms. Since the evolution time is finite and thus not perfectly adiabatic, there is a small amount of excitation throughout the process leading to a drop in overall fidelity. As the target Hamiltonian is approached $\Delta E_{gr}$ decreases considerably, making excitation of higher energy many-body states more probable, thus reducing the final fidelity. Most unwanted excitations occur during the critical region in the final step of preparation meaning that the variation rate of the Hamiltonian must be slowed to maintain fidelity with the ground state. Given the resource intensive nature of many-body time evolution we restrict ourselves to simulating the final section of the ramp as it is the most critical, testing it for different variation rates.
III.2 Introducing phase noise
Figure 4: Final fidelity (a) and total energy per particle relative to the ground state (b) for different ramp times $T_{3}$ of the final step in adiabatic state preparation described in Fig 2 for various system sizes $N$. Black line shows results for 11 sites with no phase noise, while all other results have phase noise. Optimal values of $T_{3}$ for a given $N$ which result in highest fidelity are highlighted in red. Error bars show standard error after averaging over 100 preparations with independent laser phase noise realisations. These results have been calculated using tensor network simulations using the iTensors package in Julia [38].
Figure 5: Final fidelity of prepared ground state with fixed duration $T_{3}V_{dd}=90$ and different system sizes $N$, applying a variety of scaled phase noise. (a) Final fidelity after direct scaling of applied phase noise $\phi(t)\rightarrow m\phi(t)$. (b) Final fidelity after re-scaling frequency grid of sampled spectral density by a factor of $\kappa$, moving the the noise profile peaked at $3V_{dd}\approx 2\pi\times 480$  kHz (see Fig. 3) to higher or lower frequencies. Modified spectra are renormalized to ensure that total noise power remains constant. Dashed line shows scaling for which noise is most detrimental to fidelity. Error bars show standard error after averaging 100 preparations with laser phase independent noise realizations.
Figure 6: Matrix elements of the many-body operator $\hat{\sigma}_{z}=\sum_{k}^{N}\hat{\sigma}^{k}_{z}$ for the initial Hamiltonian (a), with laser parameters $\Omega/\text{V}_{dd}=1.0$ and $\delta/\text{V}_{dd}=1.1$, as well as the final Hamiltonian (b), with parameters $\Omega/\text{V}_{dd}=0.1$ and $\delta/\text{V}_{dd}=1.1$, of the third stage of the adiabatic state preparation initialized in the ground state $|E_{0}\rangle$ for 11 sites. Dashed lines mark energies of four energy eigenstates chosen for the analysis due to their high occupancy after a time evolution of the two time independent Hamiltonians with phase noise (see appendix B Fig. 11). Points for the same color represent off-diagonal matrix elements.
When noise is added to the transverse field, the resulting dephasing of the dynamics accumulates and, over time, causes unwanted excitations. Noise is added to the system with $\phi(t)$, which is implemented as a phase term modulating $\Omega(t)$ in Eq. (II). Realizations of $\phi(t)$ are generated using an algorithm that stochastically samples a power spectrum chosen to be typical for Rydberg lasers. Fig. 3 provides the original frequency power spectrum of laser noise for a diode laser, and transforms it into a phase power spectrum that can then be used to generate unique discrete noise signals. Details on the simulation of realistic phase noise and the excitation dynamics that lead to the results in this paper are provided in the Appendices A and B.
Fig. 4 shows the final fidelity of preparing the target ground state and energy per particle for a variety of different durations of the final step of the adiabatic state preparation $T_{3}$ with added phase noise. The initial fidelity is set to 1, assuming perfect ground state preparation for steps 1 and 2. As expected, when no laser noise is present we observe constant but diminishing improvement in fidelity with increasing $T_{3}$, as the slower variation rate approaches the adiabatic limit. The addition of phase noise results in an overall drop in efficacy of the preparation, and when taken to long evolution times fidelity begins to decrease. It is important to note here that if the two initial steps in the adiabatic ramp were included in the simulation the accumulated phase noise across the full procedure would be much greater, leading to further degradation of the ground state. There is an emergent competition between diabatic and phase noise excitation, with both sources of noise exhibiting different excitation profiles and antagonistic relationships to evolution time. This leads to an emergence of an optimal $T_{3}$ at which total energy added to the system is at a minimum, corresponding to the largest achievable fidelity in state preparation. This optimal $T_{3}$ can be seen in Fig. 4 to increase with system size as the variation rate must reduce considerably to mitigate diabatic excitation irrespective of the presence of noise. This is consistent with recent analysis [39], predicting exceedingly long adiabatic ramps for larger system sizes.
Fig. 5 demonstrates the impact that laser phase noise has on the performance of the state preparation as both strength and relative frequency profile of the noise are varied for a fixed ramp duration of $T_{3}V_{dd}=90$. In the case of noise strength, Fig. 5(a) introduces a factor $m$ that linearly scales the phase noise to $\phi(t)\rightarrow m\phi(t)$. The results demonstrate the susceptibility of such adiabatic protocols to phase noise, as we observe a dramatic drop in fidelity with the target ground state in a single order of magnitude amplification. In contrast, Fig. 5(b) shows a modest improvement in fidelity as the sampled power spectrum used for noise generation is modified to have the peak of the noise moved further away from the frequencies of $\Omega$ used during the adiabatic ramp. The modification involves a linear scale factor $\kappa$ applied to the frequency grid of the original spectrum seen in the $x$-axis of Fig. 3(a), thus maintaining the relative shape of the spectrum while controlling the frequency region where the noise power is most prominent. Care was taken to ensure that the total power of the original noise spectral density is preserved, by renormalizing the modified spectrum such that its integral across the relevant frequencies is the same as that of the original. Testing a constant ramp for various modified noise spectra with peaks both above and below the original spectrum, Fig. 5 shows that moving the peak power of the phase noise away from the frequencies taken during the preparation by the optical drive $\Omega(t)$ leads to improved fidelity. Phase noise made up of frequencies that are too high or too low relative the dynamics of the system inhibit the ability of the noise to transfer power into the system, thereby causing less excitation. This suggests that even if phase noise is present in a laser, the effects of phase noise on many-body experiments can be effectively mitigated by moving dynamics far away from affected frequencies as was shown for single atoms and two qubit gates in [21, 22].
III.3 Matrix element analysis
Further insight into the behavior of noise in the initial and final stages of the adiabatic state preparation is provided through the analysis of dominant matrix elements that facilitate phase noise excitation. Given that the Hamiltonian in Eq.(II) is defined in terms of the many-body number operator $\hat{n}=\sum_{k}^{N}\hat{n}_{k}$, it is more natural to work in the rotating frame, where phase noise acting on $\Omega$ manifests itself as a fluctuating detuning $\delta$ because it alters the laser frequency seen by the atom [40, 41]. This explains why laser phase noise fluctuations on Rabi frequency can be treated as a perturbation in detuning, with a new effective detuning term $\delta(t)\rightarrow\delta_{\mathrm{eff}}(t)=\delta(t)+\dot{\phi}(t)$ acting on the system through $\hat{n}$. Fig. 6 shows the available transitions that can be induced by phase noise, plotting all matrix elements of $\hat{n}$ for the four most prominent energy eigenstates in the initial and final Hamiltonian of the adiabatic state preparation. In the case of the initial Hamiltonian in Fig. 6(a), there is adifferential writesingle dominant transition from the ground state $|E_{0}\rangle$ to the ninth excited eigenstate $|E_{9}\rangle$ at an energy of $E_{9}=1.5V_{dd}$. As $|E_{9}\rangle$ becomes excited over time, its transition elements also become relevant, and we once again see the two dominant excitation channels are to excite to $|E_{64}\rangle$ with energy $3V_{dd}$, or de-excite back to $|E_{0}\rangle$. The pattern is repeated for higher eigenstates, such that time evolving a Hamiltonian with phase noise and a constant $\Omega/V_{dd}=1.0$ leads to the formation of an excitation ladder of equally spaced peaks (see Fig. 11(a) in Appendix B). However, the picture is very different in the final Hamiltonian of the adiabatic state preparation with $\Omega/V_{dd}=0.1$ in Fig. 6(b), where we observe excitation from the ground state as well as the higher energy states strongly suppressed. For $|E_{0}\rangle$, the most probable transition has a probability that is two orders of magnitude smaller than for the original Hamiltonian, suggesting that the ground state is much more robust to excitation in the final stages of the ramp. Furthermore, the energy eigenvectors that dominated the excitation dynamics in the initial Hamiltonian appear as higher excited states, while the equidistant transition peaks of the matrix elements of the initial Hamiltonian are replaced with transitions for much smaller probabilities and clustered near their respective eigenstate. This suggests that excitation due to laser phase noise is much more constrained in the final stages of the experiment, with smaller transition rates leading to less overall excitation.
[…]
IV Thermalization
The presence of phase noise introduces energy into the system. It is then natural to ask if and how this added energy thermalizes across the many-body eigenstates. We define thermalization as a relaxation to states where the values of macroscopic observables become stationary over long periods of time, across widely differing initial conditions, and predictable by the use of statistical mechanics [29]. The mechanisms underpinning thermalization in quantum systems vary considerably from classical systems in which the process is described through chaotic dynamics in ergodic systems that lead to a Boltzmann distribution across all accessible states over long times. This understanding of thermalization is incompatible with isolated quantum systems, for which dynamical chaos and ergodicity are absent due to the linearity of time evolution and discreteness of the state-space. Instead, quantum thermalization is generally studied by comparing excitation spectra of dynamical quantum systems to a variety of thermal ensembles predicted by statistical mechanics. Such thermalization has been observed in several generic isolated quantum systems [43, 44, 45, 46, 47, 48], integrable systems that exhibit a large number of conserved quantities do not reach thermal equilibrium [49, 50, 51, 48] and instead relax to the generalized Gibbs ensemble [26, 52]. Whether or not thermalization is observed at the end of this adiabatic procedure will therefore depend strongly on the integrability of the Hamiltonian and the choice of observable.
For systems that relax to a thermal equilibrium, the underlying mechanism is most robustly described by the eigenstate thermalization hypothesis (ETH) [29]. The ETH states that for nonintegrable systems (that do not exhibit many-body localization [53, 54]) the expectation values fluctuate negligibly around eigenstates $|\Psi_{\alpha}\rangle$ that are close to the energy of the system. Thus, each individual eigenstate can be thought of as an independent thermal state. In the thermodynamic limit, the resulting state should be constant and predictable by the total energy of the system. The hypothesis also presumes that the observable of interest is ‘well behaved’, in that $\langle\Psi_{\alpha}|\hat{A}|\Psi_{\alpha}\rangle$ behaves smoothly as a function of eigenstates of the Hamiltonian $|\Psi_{\alpha}\rangle$, with no sudden discontinuities in the vicinity of the expected energy that would create a sensitivity to slight changes in relaxation dynamics. Given these conditions the ETH predicts that the long time (LT) expectation value of a few body observable $\hat{A}$ with an eigenstate $|\Psi_{\alpha}\rangle$ of a many-body system Hamiltonian $\hat{H}$ with corresponding well defined total energy $E_{\alpha}$ is equal to the thermal ensemble $\langle\hat{A}\rangle_{\text{Therm}}(E_{\alpha})$ of $\hat{A}$ at a mean energy $E_{\alpha}$
$\langle\Psi_{\alpha}|\hat{A}|\Psi_{\alpha}\rangle_{LT}=\langle\hat{A}\rangle_{\text{Therm}}(E_{\alpha}).$ (5)
Hence, the ETH provides a framework for comparing expectation values which are directly dependent on initial conditions to thermal ensembles which instead depend only on the total energy of the system. This implies a universal equilibrium for a thermalizing system across many trajectories of equivalent energy.
It will be necessary to define what is meant by a long time average $\langle\Psi_{\alpha}|\hat{A}|\Psi_{\alpha}\rangle_{LT}$ in Eq. (5). Any dynamics in a quantum system leads to correlations between excited energy eigenstates that cause fluctuations in measured expectation values as a function of time and energy difference. The time it takes for these fluctuations to die down drops with increasing system size, but for small and intermediate systems with large energy gaps this time can be too long to measure or even simulate. However, numerically a long time average can be accessed directly by considering the diagonal ensemble
$\langle\hat{A}\rangle_{LT}=\sum_{i}|c_{ii}|^{2}\langle E_{i}|\hat{A}|E_{i}\rangle=\sum_{i}|c_{ii}|^{2}\hat{A}_{ii},$ (6)
in which we sum over all eigenstates $|E_{i}\rangle$ and effectively omit correlations by not considering off-diagonal elements in the density matrix of a state. The resulting long time expectation value can then be compared to an appropriate thermal ensemble.
To understand the ETH fully it is also important to define the thermal ensemble $\langle\hat{A}\rangle_{\text{Therm}}(E_{\alpha})$. In statistical mechanics the thermal equilibrium of a dynamical system is defined in terms of the Boltzmann distribution normalized by a partition function $Z=\sum_{i}\mathrm{exp}(-\beta E_{i})$ where $\beta$ is the thermodynamic temperature of the system. An ensemble of available energy states is then scaled by the thermodynamic temperature $\beta$ which can be calculated using the total energy of the system. Adapting this to the Dirac notation, we define the canonical ensemble as
$\langle\hat{A}\rangle_{\text{canon}}=\frac{1}{Z}\sum_{i}\mathrm{exp}(-\beta E_{i})\langle E_{i}|\hat{A}|E_{i}\rangle,$ (7)
where we sum over all the energy eigenstates $|E_{i}\rangle$. Previous work has shown that for closed systems such as the one discussed in this paper the canonical ensemble is not always representative of final equilibrium [25]. In such finite systems thermalization occurs when energy in a small neighborhood of eigenstates uses the rest of the eigenstates as a thermal reservoir to dissipate into. Since the reservoir itself is not infinite, we see finite-size effects which become more prevalent at smaller system sizes. Thermalization in isolated quantum systems is typically discussed instead in terms of the microcanonical ensemble which considers only a small energy shell of energies, equally weighted and centered around the expected energy. Nevertheless, this approach is highly dependent on the density of states around the average energy of the system. The exact size of the energy shell should be small in comparison to the energy scale of the system, but large enough to include a statistically significant number of energy eigenstates. Relaxation to the microcanonical ensemble is therefore more readily observed in the middle of the energy spectrum, and for larger system sizes, where gaps between energies narrow down and density of states is higher [46]. However, the low energy spectrum which dominates in adiabatic state preparation usually exhibits a much lower density of states, leading to nonsensical microcanonical averages as very few or even no eigenstates fall within the vicinity of a given energy shell. The following section will thus evaluate the ETH using the canonical ensemble.
IV.1 Evaluating thermalization
In the case of adiabatic state preparation with experimentally relevant levels of phase noise, a comparatively small amount of energy is added to the system. This means that the dynamics will be mainly contained in the low energy spectrum where the microcanonical average is not well defined, limiting the analysis to the canonical ensemble. A value for $\beta$ of a system at a given energy can always be calculated by solving the equation
$\langle\psi|\hat{H}_{R}|\psi\rangle=\frac{1}{Z}\sum_{i}\mathrm{exp}(-\beta E_{i})\langle E_{i}|\hat{H}_{R}|E_{i}\rangle,$ (8)
where $|\psi\rangle$ represents the state of the system at the end of the adiabatic state preparation such that the left hand side of the equation gives the final energy at the end of the procedure. Once a value for $\beta$ is determined it can be used in Eq. (7) to evaluate arbitrary thermal expectation values at the appropriate energy. Given that phase noise has the potential for exciting across the full energy spectrum it is important to access the full Hilbert space, meaning that we limit ourselves to a system size of 11 sites which can be completely diagonalized.
Fig. 7 provides all the distributions required to generate long time and thermal expectation values as a function of energy relative to the ground state. Figures 7(a) and 7(b) show the two observables chosen for the comparison to be the interaction energy $\hat{H}_{int}$ as defined in Eq. (2), as well as a $\hat{O}_{Z_{2}}$ order parameter defined in Eq. (3) that measures long range $Z_{2}$ ordering that is maximally represented in the ground state. Given the relatively small Rabi frequency $\Omega=0.1V_{dd}$ at the end of the state preparation, we are investigating a Hamiltonian that is highly integrable, with structure in energy levels leading to regions of clustered eigenvalues (see Fig. 10 in Appendix B). Although neither of these observables vary smoothly as a function of energy as prescribed by the ETH, there is considerably less variance in $H_{int}$ compared to the $\hat{O}_{Z_{2}}$ order parameter meaning that small changes in excitation energy should affect expectation values of the latter more. The distributions as a function of eigenstate energy for $H_{int}$ and the $\hat{O}_{Z_{2}}$ order parameter are then weighted by the diagonal and canonical ensembles. Fig. 7(c) provides example distributions for a short preparation of $T_{3}V_{dd}=45$ and a much longer preparation at $T_{3}V_{dd}=400$ in which the Boltzmann distribution skews visibly towards higher excited states. The diagonal ensemble for the shorter $T_{3}V_{dd}=45$ shows the first excited state having the highest representation, with a majority of higher energy states left largely unexcited and thus with an lower occupation than is predicted by the Boltzmann distribution. For a much longer preparation of $T_{3}V_{dd}=400$ the first excited state drops far below the occupation of the Boltzmann distribution, as diabatic excitation is subdued. Instead, we see a general rise in occupancy across the entire energy spectrum, with a cohort of low energy eigenstates clustered below $E=1V_{dd}$ from the ground state being most prominent.
Figure 7: (a) Interaction energy $H_{int}$ defined in Eq. (2) and (b) $\hat{O}_{Z_{2}}$ ordering parameter as a function of low energy eigenstates of the target Hamiltonian with $\Omega=0.1V_{dd}$ and $\delta=1.1V_{dd}$ in a one dimensional Rydberg spin chain with 11 sites. (c) Long time ensemble (LT) $\langle\Psi|E_{i}\rangle$, and thermal ensemble (Th) as predicted by the canonical ensemble from Eq. (7), for the final step of the adiabatic state preparation with durations $\text{T}_{3}=45,400tV_{dd}$. Energy defined with respect to ground state energy.
Figure 8: Expectation values for (a) interaction energy $H_{int}$ and (b) $\hat{O}_{Z_{2}}$ ordering as a function of time taken to complete the third step in the adiabatic state preparation $T_{3}$. Thermal expectation values (green) that average over the canonical ensemble are compared to long time expectation values (orange) that average over the diagonal ensemble. Error bars show standard error over 100 preparations with laser phase noise realizations.
Averaging over the operator distributions of Figures 7(a) and 7(b) weighted by the long time and thermal ensembles of a particular $T_{3}$ result in a single expectation value. Fig. 8 provides a comparison between thermal and long time expectation values of $\hat{H}_{int}$ and the $\hat{O}_{Z_{2}}$ order parameter, across a chain of 11 sites, and for a variety of $T_{3}$. The general relationship between both types of expectation value and time $T_{3}$ is roughly consistent with the results in Fig. 4, as the optimal ramp time coincides with the lowest interaction energy in the system and therefore the highest fidelity with the $Z_{2}$ ordered ground state. For durations shorter than the optimal fidelity, where diabatic excitation dominates, we see a clear divergence between the long time and thermal expectation values. The cause of this divergence can be seen in Fig. 7(c), which shows that for the shorter preparation time of $T_{3}=45tV_{dd}$ a majority of energy eigenstates have occupation probabilities below the Boltzmann distribution. However, as the duration of $T_{3}$ passes its optimum, and laser phase noise begins to dominate excitation, we see a clear convergence of the long time and thermal expectation values. In the case of $H_{int}$ the convergence is strong as soon as $T_{3}V_{dd}=60$, coinciding with a plateau of minimum energy seen for the same system size of 11 sites in Fig. 4(b). At the optimal $T_{3}$ there is not enough energy added to the system to differentiate the two expectation values. However, this strong convergence persists as phase noise excitations heat the system. In the case of the $\hat{O}_{Z_{2}}$ order parameter, a similar divergence of the two expectation values occurs at short $T_{3}$, along with convergence around the optimal time. For longer $T_{3}$, the expectation values for this parameter behave more erratically as can be expected given the large variance of the operator as a function of energy. Despite this, there is still good convergence between the thermal and long time expectation values even for the $\hat{O}_{Z_{2}}$ operator.
Analysis of the diagonal ensemble for a ground state evolution of constant Hamiltonians at different $\Omega$ along the simulated adiabatic preparation reveals that phase noise excitation is considerably more prominent in the initial stages of the ramp (see Fig. 11 in Appendix B). When $\Omega$ is large enough to break integrability, and the phase noise is given enough time, the highly delocalized transitions that dominate excitation form energy peaks across the entire spectrum that remain visible even at the end of the simulation in Fig. 7(c). However, by the end of the state preparation when $\Omega$ decreases and the Hamiltonian becomes integrable, these excitation peaks become distorted as energy eigenstates shuffle into tightly packed clusters separated by large energy gaps. There is an overall drop in transition probability as highly delocalized excitations disappear, limiting the energy mobility around the system to the local heating of neighboring eigenstates predicted by Fig. 6(b) and shown in Fig. 11(c). The combination of aggressive delocalized excitation that facilitates rapid energy transfer in the first half of the ramp, and the slow localized heating of this transferred energy in the latter half seems to excite the system in a way such that averaged behavior is lockstep with the uniform predictions of the Boltzmann distribution, albeit the distributions differ considerably. In this way, we can say the adiabatic state preparation facilitates the convergence of the long-time and thermal ensemble. However, when phase noise is left uncontrolled in a non-integrable regime it leads to a diagonal ensemble that rises above the canonical ensemble and does not resemble thermal equilibrium. Similarly, evolving in a Hamiltonian that is fully or close to integrable leads to much less total energy transfer, with strictly localized excitation inhibiting the uniform excitation of the full energy spectrum.
V Conclusion
This study has examined the impact of laser phase noise on adiabatic state preparation in a one-dimensional Rydberg spin chain, focusing on its role in driving excitations and influencing thermalization dynamics. By incorporating phase noise sampled from realistic laser power spectra, the simulations provide a realistic assessment of how coherence errors in the transverse field affect quantum state preparation. The results highlight a competition between diabatic and noise-induced excitations, leading to an optimal ramp duration that minimizes total energy absorption and improves final fidelities of the prepared state. The role of noise resonance and noise scaling is also examined, showing that phase noise is most disruptive when its dominant frequency components are resonant with key system excitations. Adjusting the spectral profile of the noise or shifting the system’s operational frequency away from the most affected regions improves state preparation fidelity. Additionally, the results demonstrate that increasing the overall noise strength leads to a rapid decline in performance, underscoring the need for noise suppression techniques in scalable neutral atom platforms.
Excitation due to phase noise is found to depend strongly on the integrability of the system, with non-integrable regimes seen in the simulated adiabatic ramp facilitating energy delocalization, while integrable regimes lead to energy clustering and large spectral gaps which limit the energy transfer to low probability localized excitations. The interplay of these two regimes has direct implications on how energy is dissipated across the system, as we investigate whether this resembles many-body thermalization. Analysis of long-time expectation values indicates that noise-driven excitations can, under certain conditions, lead to thermalization consistent with the eigenstate thermalization hypothesis (ETH). The observed convergence between diagonal and thermal ensembles suggests that in these systems, the excitations generate correlation functions that are close to what would be expected in a thermal ensemble. However, we should note that in important parts of the parameter space, these systems are approximately integrable and will generally lead to deviations from thermalization – even more so when the noise strength is high enough to induce excessive heating. In contrast, in strictly integrable regimes, the absence of delocalized excitation channels restricts energy transfer across the spectrum and for higher energy states in particular, also preventing the system from reproducing thermal predictions.
Our results have broad implications for quantum simulation and computation with Rydberg atom arrays, especially as system sizes increase and noise mitigation becomes a growing challenge. Future research in this topic could explore a broader range of phase noise spectra [24], higher-dimensional Rydberg geometries, different Rydberg accessible Hamiltonians, as well as observing the effects of phase noise on different interacting Hamiltonians. Also of interest is a parallel investigation of phase noise in the Van der Waals Rydberg interaction regime in which interactions scale as $1/r^{6}$ as opposed to the $1/r^{3}$ scaling of dipole-dipole interactions, and has become common in experimental setups in recent years [55, 56, 57, 58].
Aknowledgements
We are grateful to Sebastian Schmidt, Sridevi Kuriyattil, and the members of the SQUARE project team for helpful discussions. This work is supported by the EPSRC (Grant No. EP/T005386/1) and M Squared Lasers Ltd. The data presented in this work are available at [59].
Appendix A Noise Generation
The method used for noise generation is based on the TK95 algorithm [60] for the generation of independent time signals from a frequency power spectrum by applying a complex Gaussian filter followed by the inverse Fourier transform. To begin with, a sample spectral density of frequency noise fluctuations $S_{\nu}$ is chosen (see Fig. 3).d The frequency spectral density can be transformed into a phase spectral density by applying $S_{\nu}=\nu^{2}S_{\phi}$. In case a voltage power spectral density is provided $S_{V}$, we refer the reader to [61] where the conversion to $S_{\phi}$ is discussed in detail. To create a real noise signal, a double-sided conjugate symmetric spectrum must be generated. Given that the provided spectrum includes only positive frequencies, it is necessary to sample the spectrum twice using a complex Gaussian factor that is conjugated for negative frequencies. A frequency spectrum $S_{n}$ sampled from the phase noise power spectrum $S_{\phi}$ is defined by
$S_{n}\bigg{[}\frac{\mathrm{rad}^{2}}{\mathrm{MHz}}\bigg{]}^{\frac{1}{2}}$ $=\mathcal{N}(0)\sqrt{\frac{S_{\phi}}{2}}+i\mathcal{N}(0)\sqrt{\frac{S_{\phi}}{2}},$ (9)
$S_{n}(-\nu)$ $=S_{n}^{*}(\nu)\qquad k\in[0,\mathrm{\nu}_{Nyq}].$
where the relevant units are shown in square brackets and $\mathcal{N}(0)$ is a Gaussian random number used to sample $S_{\phi}$ at each frequency ±$\pm k$ up until the Nyquist frequency $\nu_{Nyq}$. Once a trajectory of a phase noise spectrum $S_{n}$ is generated, the corresponding phase noise signal $\phi(t)$ is derived by an inverse Fourier transform
$\phi(t)=N\frac{\sqrt{2\Delta\nu}}{2\pi}\sum^{N-1}_{\nu=0}S_{n}(\nu)e^{2\pi i\nu t/N}.$ (10)
with the addition of a normalization factor where $N$ is the total number of frequency bins $k$ in $S_{w}$ (positive and negative) and a factor of $(2\pi)^{-1}$ is added to account for the Fourier transform in frequency space as opposed to angular frequency. The result of this algorithm is thus a time signal of phase noise $\phi(t)$ in angular units that can then be applied directly to the time evolution of the many-body Hamiltonian in Eq. (II). To test the validity of the generated noise, we compare the autocorrelation function taken directly from the sampled spectrum $A_{S_{\phi}}(\tau)$ with the averaged autocorrelation across 100 generated noise signals $A_{\phi}$. For this we use the fact that the autocorrelation function can be attained via direct Fourier transform of a given spectral density. Given that $S_{\phi}$ is provided as a discrete series in frequency space, approximating the inverse Fourier transform to a Riemann sum gives
Figure 9: Comparison of the autocorrelation function of the original power density spectrum $A_{S_{\phi}}$ (blue line) and an average of 100 autocorrelations corresponding to independently generated noise signals $A_{\phi}$ (pink line). Error bars represent standard error and show good agreement in average noise behavior.
$A_{S_{\phi}}(\tau)$ $=\frac{N\Delta\nu}{2\pi}\sum_{\nu}S_{\phi}(\nu)e^{2\pi i\nu\tau/N},$ (11)
where $A_{S_{\phi}}(\tau)$ is the autocorrelation function of the noise at a given time lag $\tau$. In the case of atocorrelations for the generated phase noise, the autocorrelation of a given time series can be calculated using
$A_{\phi}(\tau)=$ $\langle\phi_{0}\phi_{\tau}\rangle=\frac{E[(\phi_{0}-E[\phi_{0}])(\phi_{\tau}-E[\phi_{\tau}])]}{\sigma(\phi_{\tau})\sigma(\phi_{0})},$ (12)
where $\phi_{0}$ and $\phi_{\tau}$ are the same noise signal starting from $t=0$ and a lag of $t=\tau$ respectively, and $\sigma$ gives the standard deviation. Note that the numerator here is just the covariance between $\phi_{0}$ and $\phi_{\tau}$ where $E$ signifies the expected value of the given time series. Fig. 9 shows the agreement in autocorrelation functions of the original power spectrum $S_{\phi}$ and an average of 100 independently generated noise realizations.
Appendix B Excitation dynamics
Having discussed phase noise generation, we now demonstrate some of the features in the excitation dynamics in our system. Consider again the Hamiltonian in Eq. II where a given noise array $\phi(t)$ appears as a small phase modulation on the Rabi frequency of the laser drive at every time step in the evolution. The following section will first highlight the reflection symmetry that governs the dynamics in this Hamiltonian, as well as exploring the different features in the diagonal ensemble after phase noise excitation in a time independent Hamiltonian.
Figure 10: A plot of symmetric energy eigenvalues of the Rydberg Hamiltonian defined in Eq. (II) as a function of Rabi frequency $\Omega$ for a constant detuning of $\delta=1.1V_{dd}$.Increasing $\Omega$ strengthens the transverse field, leading to a rapid breakdown in integrability as initially clustered eigenstates spread out and merge.
B.1 Reflection symmetry
Importantly, this Hamiltonian has a reflection symmetry in the computational basis. Given that we limit ourselves to odd numbers of sites, a reflection operator that flips all sites about the center of the chain can be defined in terms of $\sigma_{+}=(\sigma_{x}+i\sigma_{y})/2$ and $\sigma_{-}=(\sigma_{x}-i\sigma_{y})/2$ operators as
$\hat{R}=\prod_{i=1}^{N/2}\bigg{(}$ $\hat{\sigma}_{+}^{(i)}\hat{\sigma}_{-}^{(N-i+1)}+\hat{\sigma}_{-}^{(i)}\hat{\sigma}_{+}^{(N-i+1)}$ (13)
$+\frac{1}{2}\big{(}1+\hat{\sigma}_{z}^{(i)}\hat{\sigma}_{z}^{(N-i+1)}\big{)}\bigg{)}.$
Figure 11: Above are plots of the diagonal ensemble as a function of energy eigenstates after a time evolution with four time independent Hamiltonians and laser phase noise. The ground state of the Hamiltonian from Eq. (II) is evolved with noise for a total time of $tV_{dd}=400$, with constant parameters $V_{dd}=1/2\pi$MHz, $\delta/V_{dd}=1.1$, and (a) $\Omega/V_{dd}=1.0$, (b) $\Omega/V_{dd}=0.6$, (c) $\Omega/V_{dd}=0.4$, (d) $\Omega/V_{dd}=0.2$.The plots are reproduced for two different noise scale factors $\kappa$ which move the peak power of the sampled noise PSD to higher and lower frequencies in Fourier space (see Fig. 5). In orange we have the diagonal ensemble for $\kappa=0.9$ which was shown to cause most excitation during the adiabatic state preparation, while in blue we have $\kappa=0.5$ corresponding to a PSD that is further from resonance with the system causing less excitation. Error bars plot standard error across 100 time evolutions with independent laser phase noise signals, but are too small to be seen suggesting consistent excitation pattern.
The first two terms in the operator flip sites at either end of the chain while the third term ensures the preservation of the spin alignment if both sites are in the same state. The Hamiltonian can be shown to commute with this reflection operator, leading to two independent symmetry sectors of eigenstates made up of symmetric and anti-symmetric superpositions of computational states. For open boundary one-dimensional spin chains with an odd number of sites, there will always be computational states that are invariant under reflection. For example, the crystalline states discussed in Fig. 1 that dominate the ground state for low values of $\Omega$. Such reflection invariant computational states are only represented in the symmetric energy eigenstates. Furthermore, it can be shown that diabatic excitation, governed by changes to $\Omega$, as well as phase noise excitation, governed by changes in $\phi$, both respect this symmetry. This means that if all energy is initially in one of these symmetry sectors as is the case in the symmetric ground state at the beginning of the adiabatic state preparation, all dynamics will be confined to that symmetry sector. This effectively inhibits excitation in approximately half of the Hilbert space, thus halving the density of states that the noise can dissipate into.
Fig. 10 shows the separation of eigenenergies corresponding to symmetric eigenstates as a function of Rabi frequency $\Omega/V_{dd}$ and with constant detuning $\delta=1.1V_{dd}$, which are relevant parameters for the third step in the simulated adiabatic state preparation. Tracing the bottom lone which represents the ground state we see the minimum ground state energy gap corresponding to the critical region of the preparation that occurs between $\Omega=0.2-0.3V_{dd}$. Furthermore, the total energy span of the Hilbert space increases from $10V_{dd}$ to $6V_{dd}$ as $\Omega/V_{dd}$ increases. Initially, clusters of energy eigenvalues form an underlying structure of clustered energy eigenstates that is consistent with what one might expect in an Ising Hamiltonian, but as $\Omega/V_{dd}$ we see this structure break down as the clusters merge. This merging corresponds to the breaking of integrability that occurs when a transverse field is applied to the Ising Hamiltonian, which has direct implications for phase noise excitation.
B.2 Laser phase noise excitations in time independent Hamiltonians
While it is simple to treat the adiabatic state preparation as a black box and analyze only the final state, getting a deeper understanding of the complicated excitation dynamics of laser phase noise a time dependent Hamiltonian with ever changing energy level structure is much harder. Fig. 11 provides a small glimpse by choosing four different Hamiltonians at various values of $\Omega$ that occur throughout the simulated adiabatic state preparation, performing a time independent evolution of the ground state with laser phase noise. The resulting diagonal ensembles for two separate noise signals generated with different power profiles demonstrate the diminishing effects of laser phase noise as $\Omega$ is reduced and the system becomes increasingly integrable. The evolution of the ground state in the initial Hamiltonian of the preparation of the state in Fig. 11(a) results in the formation of well-defined equidistant peaks as the energy of the phase noise excitation cascades from one peak to the next. This excitation profile follows the matrix element predictions of Fig. 6(a) which plot the transition spectra of energy eigenstates corresponding to the four largest peaks, for which excitation dynamics is dominated by single delocalized transitions that excite the entire energy spectrum. Furthermore, although the excitation patterns for the two noise signals used in Fig. 11(a) are identical, the noise signal corresponding to a PSD rescaled by $\kappa=0.9$ known to result in the worst performance in state preparation (see Fig. 5(b)) results in a final state that has ground state fidelity of $\langle\psi_{gr}|\psi\rangle_{\kappa=0.9}=0.58$, compared to $\langle\psi_{gr}|\psi\rangle_{\kappa=0.5}=0.77$ for $\kappa=0.5$. In other words, the amount of phase noise excitation is highly dependent on the level of power a given realization has in the frequency region of the dominant transitions of the Hamiltonian. The Hamiltonian in Fig. 11(b) has $\Omega/V_{dd}=0.6$ which occurs near the middle of the simulated state preparation, and despite a contraction of the energy spectrum we still observe peaked features as for the initial Hamiltonian. This suggests that although the exact energy eigenstates might differ, the excitation dynamics in the initial stages of the state preparation exhibit similar behavior. However, upon closer inspection we see that the ground state fidelities have risen to $\langle\psi_{gr}|\psi\rangle_{\kappa=0.9}=0.67$ and $\langle\psi_{gr}|\psi\rangle_{\kappa=0.5}=0.81$, suggesting that the approach towards integrability is already affecting the total levels of excitation. Fig. 11(c) shows evolution with a Hamiltonian with $\Omega/V_{dd}=0.6$. We see a clear degradation of the peaked excitation spectra seen in the previous two plots, as the excitation dynamics shift towards more localized transition that lead to less overall energy mobility as the highest energy states are left largely unexcited. A drop in overall excitation is also evidenced by the rise in final ground state fidelities of $\langle\psi_{gr}|\psi\rangle_{\kappa=0.9}=0.84$ and $\langle\psi_{gr}|\psi\rangle_{\kappa=0.5}=0.82$, as we see that the low-frequency noise signal generated using $\kappa=0.5$ is now marginally more likely to excite the system. The transition to localized excitation is completed by the final stages of the Hamiltonian as seen in Figure Fig. 11(d) where evolution with a transverse field of $\Omega/V_{dd}=0.6$ leads to much lower levels of excitation as final ground state fidelities rise further to $\langle\psi_{gr}|\psi\rangle_{\kappa=0.9}=0.96$ and $\langle\psi_{gr}|\psi\rangle_{\kappa=0.5}=0.95$. Apart from the excitation out of the ground state itself, all phase noise excitation is localized to a small energy region surrounding the state with an exponential drop off in excitation as you get further away.